I can't figure out how to do this problem.
The sum of three times a number and 7 more than the number is the same as the difference between-11 and twice the number. What is the number?

Answer by stanbon(75887)

The sum of three times a number and 7 more than the number is the same as the difference between-11 and twice the number. What is the number?
----
3x + (3x+7) = -11-2x
-------
10x+7 = -11-2x
12x = -18
x = -3/2
==============
Cheers,
Stan H.
==============